# The Exact Relativistic Scalar Quasibound States of The Dyonic Kerr-Sen Black Hole: Quantized Energy, and Hawking Radiation

### v1

@@ Abstract
We consider Klein-Gordon equation in the Dyonic Kerr-Sen black hole background, which is the charged rotating axially symmetric solution of the Einstein-Maxwell-Dilaton-Axion theory of gravity. The black hole incorporates electric, magnetic, dilatonic and axionic charges and is constructed in 3+1 dimensional spacetime. We begin our investigations with the construction of the scalar field’s governing equation, i.e., the covariant Klein-Gordon equation. With the help of the ansatz of separation of variables, we successfully separate the polar part, and find the exact solution in terms of Spheroidal Harmonics, while the radial exact solution is obtained in terms of the Confluent Heun function. The quantization of the quasibound state is done by applying the polynomial condition of the Confluent Heun function that gives rise to discrete complex-valued energy levels for massive scalar fields. The real part is the scalar field relativistic quantized energy, while the imaginary part represents the quasibound states’s decay. We present all of the sixteen possible exact energy solutions for both massive and massless scalars. We also present the investigation the Hawking radiation of the Dyonic Kerr-Sen black hole’s apparent horizon, via the Sigurd-Sannan method by making use of the obtained exact scalar wave functions. The radiation distribution function, and the Hawking temperature are successfully obtained.
@@ I Introduction
After the gravitational wave signal of a binary black hole merger was finally detected for the first time on 14 September 2015 205 , quite recently, the Hawking radiation of an optical black hole analog drori was also finally observed. This makes the investigation of the black hole spectroscopy a new emerging field of major physical interest. The quasibound states, quasinormal modes, and shadows of black holes are among the most interesting characteristics of such astrophysical objects, and they appear in the observationally measurable spectra that are released as the particles cross into the black hole 100 .
@@ III.1 The Klein-Gordon equation
$-{\hbar}^{2}\nabla_{\mu}\nabla^{\mu}\psi+m^{2}c^{2}=0,$ (55)
@@
$\psi\left(t,r,\theta,\phi\right)=e^{i\frac{E}{c}ct-im_{l}\phi}R\left(r\right)T\left(\theta\right).$ (64)
@@ III.3 The Radial Wave Equation
$F.T.=\frac{1}{\delta_{r}}\frac{1}{y\left(y+1\right)}\left\{\frac{\Omega}{r_{s}}\left[\delta^{2}_{r}y^{2}+2y\delta_{r}(r_{+}-2d)\right]\right.\\ \left.+\underbrace{\frac{\Omega}{r_{s}}\left[r_{+}\left(r_{+}-2d\right)-k^{2}+a^{2}\right]-m_{l}a}_{K_{1}}\right\}.$ (83)
@@
$F.T.=\frac{\Omega}{r_{s}}\delta_{r}+\frac{1}{y}\frac{K_{1}}{\delta_{r}}\\ +\frac{1}{y+1}\left\{\underbrace{\frac{\Omega}{r_{s}}\left(r_{+}+r_{-}{-4d}\right)-\frac{K_{1}}{\delta_{r}}}_{{K_{3}}}\right\}.$ (84)
@@
$S.T.=-\frac{1}{y\left(y+1\right)}\left\{\frac{\Omega^{2}_{0}}{r^{2}_{s}}\left(\delta^{2}_{r}y^{2}+2\delta_{r}y\left(r_{+}-2d\right)\right)\ \right.\\ \left.+\underbrace{\frac{\Omega^{2}_{0}}{r^{2}_{s}}\left(r_{+}\left(r_{+}-2d\right)-k^{2}+a^{2}\right)+K^{m_{l}}_{l}}_{K_{2}}\right\}.$ (86)
@@
$S.T.=-\frac{\Omega^{2}_{0}}{r^{2}_{s}}\delta^{2}_{r}-\frac{K_{2}}{y}\\ +\frac{1}{y+1}\left[\underbrace{\frac{\Omega^{2}_{0}}{r^{2}_{s}}\delta^{2}_{r}-2\delta_{r}(r_{+}-2d)\frac{\Omega^{2}_{0}}{r^{2}_{s}}+K_{2}}_{K_{4}}\right].$ (87)
@@
$\gamma=\pm\frac{2i}{\delta_{r}}\left[\frac{\Omega}{r_{s}}\left(r_{-}\left(r_{-}-2d\right)-k^{2}+a^{2}+2d\delta_{r}\right)-m_{l}a\right],$ (101)
$\delta=-\frac{\delta_{r}}{r^{2}_{s}}\left(r_{+}+r_{-}-4d\right)\left[2\Omega^{2}-\Omega^{2}_{0}\right],$ (102)
$\eta=-\frac{2}{\delta^{2}_{r}}\left(\frac{\Omega}{r_{s}}\left(r_{+}\left(r_{+}-2d\right)-k^{2}+a^{2}\right)-m_{l}a\right)\times\\ \left(\frac{\Omega}{r_{s}}\left(-r_{+}\left(r_{+}-2d\right)+2r_{-}\left(r_{+}-2d\right)-k^{2}+a^{2}\right)-m_{l}a\right)\\ -\frac{\Omega^{2}_{0}}{r^{2}_{s}}\left(r_{+}\left(r_{+}-2d\right)-k^{2}+a^{2}\ \right)-K^{m_{l}}_{l}.$ (103)
@@
$\psi=e^{i{\frac{E}{\hbar c}}ct}S_{l}^{m_{l}}\left(\theta,\phi\right)e^{-\frac{1}{2}\alpha\left(\frac{r-r_{+}}{\delta_{r}}\right)}{\left(\frac{r-r_{-}}{\delta_{r}}\right)}^{\frac{1}{2}\gamma}\\ \times\left[A{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{\frac{1}{2}\beta}\operatorname{HeunC}\left(\alpha,\beta,\gamma,\delta,\eta,-\frac{r-r_{+}}{\delta_{r}}\right)\right.\\ \left.+B{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{-\frac{1}{2}\beta}\operatorname{HeunC}\left(\alpha,-\beta,\gamma,\delta,\eta,-\frac{r-r_{+}}{\delta_{r}}\right)\right],$ (104)
@@ III.4.1 The Quasistationary Modes
$-\frac{{\left(r_{+}+r_{-}-4d\right)}\left[2\Omega^{2}-\Omega^{2}_{0}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}+\frac{i}{\delta_{r}}\left\{\frac{\Omega}{r_{s}}\left(r_{+}^{2}\right.\right.\\ \left.\left.+r_{-}\left(r_{-}-4d\right)+2\left(a^{2}-k^{2}\right)\right)-2m_{l}a\right\}=-n.$ (107)
@@ 2. For the case with $\alpha_{+},\beta_{+},\gamma_{-}$, we obtain,
$-\frac{{\left(r_{+}+r_{-}-4d\right)}\left[-\Omega^{2}_{0}+2\Omega^{2}+2i\Omega\sqrt{\Omega^{2}_{0}-\Omega^{2}}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}=-n.$ (108)
@@ 3. For the case with $\alpha_{+},\beta_{-},\gamma_{+}$ , we obtain,
$-\frac{{\left(r_{+}+r_{-}-4d\right)}\left[-\Omega^{2}_{0}+2\Omega^{2}-2i\Omega\sqrt{\Omega^{2}_{0}-\Omega^{2}}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}=-n.$ (109)
@@ 4. For the case with $\alpha_{+},\beta_{-},\gamma_{-}$, we obtain,
$-\frac{{\left(r_{+}+r_{-}-4d\right)}\left[2\Omega^{2}-\Omega^{2}_{0}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}-\frac{i}{\delta_{r}}\left\{\frac{\Omega}{r_{s}}\left(r_{+}^{2}\right.\right.\\ \left.\left.+r_{-}\left(r_{-}-4d\right)+2\left(a^{2}-k^{2}\right)\right)-2m_{l}a\right\}=-n.$ (110)
@@ 5. For the case with $\alpha_{-},\beta_{+},\gamma_{+}$, we obtain,
$\frac{{\left(r_{+}+r_{-}-4d\right)}\left[2\Omega^{2}-\Omega^{2}_{0}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}+\frac{i}{\delta_{r}}\left\{\frac{\Omega}{r_{s}}\left(r_{+}^{2}\right.\right.\\ \left.\left.+r_{-}\left(r_{-}-4d\right)+2\left(a^{2}-k^{2}\right)\right)-2m_{l}a\right\}=-n.$ (111)
@@ 6. For the case with $\alpha_{-},\beta_{+},\gamma_{-}$ , we obtain,
$\frac{{\left(r_{+}+r_{-}-4d\right)}\left[-\Omega^{2}_{0}+2\Omega^{2}-2i\Omega\sqrt{\Omega^{2}_{0}-\Omega^{2}}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}=-n.$ (112)
@@ 7. For the case with $\alpha_{-},\beta_{-},\gamma_{+}$ , we obtain,
$\frac{{\left(r_{+}+r_{-}-4d\right)}\left[-\Omega^{2}_{0}+2\Omega^{2}+2i\Omega\sqrt{\Omega^{2}_{0}-\Omega^{2}}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}=-n.$ (113)
@@ 8. For the case with $\alpha_{-},\beta_{-},\gamma_{-}$, we obtain,
$\frac{{\left(r_{+}+r_{-}-4d\right)}\left[2\Omega^{2}-\Omega^{2}_{0}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}-\frac{i}{\delta_{r}}\left\{\frac{\Omega}{r_{s}}\left(r_{+}^{2}\right.\right.\\ \left.\left.+r_{-}\left(r_{-}-4d\right)+2\left(a^{2}-k^{2}\right)\right)-2m_{l}a\right\}=-n.$ (114)
Note that the scalar’s energy levels in a rotating black hole spacetime possess hyperfine splitting, where the energy depends on the azimuthal quantum number $m_{l}$ and is directly coupled to the black hole’s spin parameter $a$. This is analogous to the Zeeman effect, which occurs when a hydrogen atom is immersed in a magnetic environment. The existence of the term can be understood as an interaction between the orbiting scalar field, with magnetic state $m_{l}$, and the black hole’s angular momentum $a$.
Now, let us investigate the energy levels in the limit $Er_{s}\to 0$. For modes with $\beta_{+},\gamma_{+}$ and $\beta_{-},\gamma_{-}$, we obtain,
$E_{n}=E_{0}\sqrt{1-{\left[\frac{\frac{E_{0}r_{s}}{\hbar c}\frac{{\left(r_{+}+r_{-}-4d\right)}}{r_{s}}}{2\left(i\frac{2m_{l}a}{\delta_{r}}+n\right)}\right]}^{2}},$ (115)
$E_{n}-E_{0}\approx-\frac{E_{0}}{2}{\left[\frac{\frac{E_{0}r_{s}}{\hbar c}\frac{{\left(r_{+}+r_{-}-4d\right)}}{r_{s}}}{2\left(i\frac{2m_{l}a}{\delta_{r}}+n\right)}\right]}^{2}.$ (116)
@@
$E_{n}=E_{0}\sqrt{1-{\left[\frac{\frac{E_{0}r_{s}}{\hbar c}\frac{{\left(r_{+}+r_{-}-4d\right)}}{r_{s}}}{2n}\right]}^{2}},$ (117)
$E_{n}-E_{0}\approx-\frac{E_{0}}{2}{\left[\frac{\frac{E_{0}r_{s}}{\hbar c}{\frac{{\left(r_{+}+r_{-}-4d\right)}}{r_{s}}}}{2n}\right]}^{2}.$ (118)
@@ 1. For the case with $\alpha_{+},\beta_{+},\gamma_{+}$, we obtain,
$\Omega_{2}=\frac{r_{s}\left(in(r_{+}-r_{-})+2m_{l}a\right)}{2\left(r_{-}(r_{-}-4d)+2r_{+}d+a^{2}-k^{2}\right)}.$ (120)

### v2

@@ Abstract
We consider Klein-Gordon equation in the Dyonic Kerr-Sen black hole background, which is the charged rotating axially symmetric solution of the Einstein-Maxwell-Dilaton-Axion theory of gravity. The black hole incorporates electric, magnetic, dilatonic and axionic charges and is constructed in 3+1 dimensional spacetime. We begin our investigations with the construction of the scalar field’s governing equation, i.e., the covariant Klein-Gordon equation. With the help of the ansatz of separation of variables, we successfully separate the polar part, and find the exact solution in terms of Spheroidal Harmonics, while the radial exact solution is obtained in terms of the Confluent Heun function. The quantization of the quasibound state is done by applying the polynomial condition of the Confluent Heun function that gives rise to discrete complex-valued energy levels for massive scalar fields. The real part is the scalar field relativistic quantized energy, while the imaginary part represents the quasibound states’s decay. We present all of the sixteen possible exact energy solutions for both massive and massless scalars. We also present the investigation the Hawking radiation of the Dyonic Kerr-Sen black hole’s apparent horizon, via the Sigurd-Sannan method by making use of the obtained exact scalar wave functions. The radiation distribution function, and the Hawking temperature are also obtained.
@@ I Introduction
After the gravitational wave signal of a binary black hole merger was finally detected for the first time on 14 September 2015 205 , quite recently, the Hawking radiation of an optical black hole analog drori was also finally observed. This makes the investigation of the black hole spectroscopy a new emerging field of major physical interest. The quasibound states (QBS), quasinormal modes (QNMs), and shadows of black holes are among the most interesting characteristics of such astrophysical objects, and they appear in the observationally measurable spectra that are released as the particles cross into the black hole 100 .
@@ III.1 The Klein-Gordon equation
$-{\hbar}^{2}\nabla_{\mu}\nabla^{\mu}\psi+m^{2}c^{2}\psi=0,$ (55)
@@
$\psi\left(t,r,\theta,\phi\right)=e^{-i\frac{E}{c}ct+im_{l}\phi}R\left(r\right)T\left(\theta\right).$ (64)
@@ III.3 The Radial Wave Equation
$F.T.=\frac{1}{\delta_{r}}\frac{1}{y\left(y+1\right)}\left\{\frac{\Omega}{r_{s}}\left[\delta^{2}_{r}y^{2}+2y\delta_{r}(r_{+}-d)\right]\right.\\ \left.+\underbrace{\frac{\Omega}{r_{s}}\left[r_{+}\left(r_{+}-2d\right)-k^{2}+a^{2}\right]-m_{l}a}_{K_{1}}\right\}.$ (83)
@@
$F.T.=\frac{\Omega}{r_{s}}\delta_{r}+\frac{1}{y}\frac{K_{1}}{\delta_{r}}\\ +\frac{1}{y+1}\left\{\underbrace{\frac{\Omega}{r_{s}}\left(r_{+}+r_{-}-2d\right)-\frac{K_{1}}{\delta_{r}}}_{{K_{3}}}\right\}.$ (84)
@@
$S.T.=-\frac{1}{y\left(y+1\right)}\left\{\frac{\Omega^{2}_{0}}{r^{2}_{s}}\left(\delta^{2}_{r}y^{2}+2\delta_{r}y\left(r_{+}-d\right)\right)\ \right.\\ \left.+\underbrace{\frac{\Omega^{2}_{0}}{r^{2}_{s}}\left(r_{+}\left(r_{+}-2d\right)-k^{2}+a^{2}\right)+K^{m_{l}}_{l}}_{K_{2}}\right\}.$ (86)
@@
$S.T.=-\frac{\Omega^{2}_{0}}{r^{2}_{s}}\delta^{2}_{r}-\frac{K_{2}}{y}\\ +\frac{1}{y+1}\left[\underbrace{\frac{\Omega^{2}_{0}}{r^{2}_{s}}\delta^{2}_{r}-2\delta_{r}(r_{+}-d)\frac{\Omega^{2}_{0}}{r^{2}_{s}}+K_{2}}_{K_{4}}\right].$ (87)
@@
${\gamma}=\pm\frac{2i}{\delta_{r}}\left[\frac{\Omega}{r_{s}}\left(r_{-}\left(r_{-}-2d\right)-k^{2}+a^{2}\right)-m_{l}a\right],$ (101)
$\delta=-\frac{\delta_{r}}{r^{2}_{s}}\left(r_{+}+r_{-}-2d\right)\left[2\Omega^{2}-\Omega^{2}_{0}\right],$ (102)
$\eta=-\frac{2}{\delta^{2}_{r}}\left(\frac{\Omega}{r_{s}}\left(r_{+}\left(r_{+}-2d\right)-k^{2}+a^{2}\right)-m_{l}a\right)\times\\ \left(\frac{\Omega}{r_{s}}\left(-r_{+}^{2}+2r_{-}\left(r_{+}-d\right)-k^{2}+a^{2}\right)-m_{l}a\right)\\ -\frac{\Omega^{2}_{0}}{r^{2}_{s}}\left(r_{+}\left(r_{+}-2d\right)-k^{2}+a^{2}\ \right)-K^{m_{l}}_{l}.$ (103)
@@
$\psi=e^{-i{\frac{E}{\hbar c}}ct}S_{l}^{m_{l}}\left(\theta,\phi\right)e^{-\frac{1}{2}\alpha\left(\frac{r-r_{+}}{\delta_{r}}\right)}{\left(\frac{r-r_{-}}{\delta_{r}}\right)}^{\frac{1}{2}\gamma}\\ \times\left[A{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{\frac{1}{2}\beta}\operatorname{HeunC}\left(\alpha,\beta,\gamma,\delta,\eta,-\frac{r-r_{+}}{\delta_{r}}\right)\right.\\ \left.+B{\left(\frac{r-r_{+}}{\delta_{r}}\right)}^{-\frac{1}{2}\beta}\operatorname{HeunC}\left(\alpha,-\beta,\gamma,\delta,\eta,-\frac{r-r_{+}}{\delta_{r}}\right)\right],$ (104)
@@ III.4.1 The Quasistationary Modes
$-\frac{{\left(r_{+}+r_{-}-2d\right)}\left[2\Omega^{2}-\Omega^{2}_{0}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}+\frac{i}{\delta_{r}}\left\{\frac{\Omega}{r_{s}}\left(r_{+}(r_{+}-2d)\right.\right.\\ \left.\left.+r_{-}\left(r_{-}-2d\right)+2\left(a^{2}-k^{2}\right)\right)-2m_{l}a\right\}=-n.$ (107)
@@ 2. For the case with $\alpha_{+},\beta_{+},\gamma_{-}$, we obtain,
$-\frac{{\left(r_{+}+r_{-}-2d\right)}\left[-\Omega^{2}_{0}+2\Omega^{2}+2i\Omega\sqrt{\Omega^{2}_{0}-\Omega^{2}}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}=-n.$ (108)
@@ 3. For the case with $\alpha_{+},\beta_{-},\gamma_{+}$ , we obtain,
$-\frac{{\left(r_{+}+r_{-}-2d\right)}\left[-\Omega^{2}_{0}+2\Omega^{2}-2i\Omega\sqrt{\Omega^{2}_{0}-\Omega^{2}}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}=-n.$ (109)
@@ 4. For the case with $\alpha_{+},\beta_{-},\gamma_{-}$, we obtain,
$-\frac{{\left(r_{+}+r_{-}-2d\right)}\left[2\Omega^{2}-\Omega^{2}_{0}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}-\frac{i}{\delta_{r}}\left\{\frac{\Omega}{r_{s}}\left(r_{+}(r_{+}-2d)\right.\right.\\ \left.\left.+r_{-}\left(r_{-}-2d\right)+2\left(a^{2}-k^{2}\right)\right)-2m_{l}a\right\}=-n.$ (110)
@@ 5. For the case with $\alpha_{-},\beta_{+},\gamma_{+}$, we obtain,
$\frac{{\left(r_{+}+r_{-}-2d\right)}\left[2\Omega^{2}-\Omega^{2}_{0}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}+\frac{i}{\delta_{r}}\left\{\frac{\Omega}{r_{s}}\left(r_{+}(r_{+}-2d)\right.\right.\\ \left.\left.+r_{-}\left(r_{-}-2d\right)+2\left(a^{2}-k^{2}\right)\right)-2m_{l}a\right\}=-n.$ (111)
@@ 6. For the case with $\alpha_{-},\beta_{+},\gamma_{-}$ , we obtain,
$\frac{{\left(r_{+}+r_{-}-2d\right)}\left[-\Omega^{2}_{0}+2\Omega^{2}-2i\Omega\sqrt{\Omega^{2}_{0}-\Omega^{2}}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}=-n.$ (112)
@@ 7. For the case with $\alpha_{-},\beta_{-},\gamma_{+}$ , we obtain,
$\frac{{\left(r_{+}+r_{-}-2d\right)}\left[-\Omega^{2}_{0}+2\Omega^{2}+2i\Omega\sqrt{\Omega^{2}_{0}-\Omega^{2}}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}=-n.$ (113)
@@ 8. For the case with $\alpha_{-},\beta_{-},\gamma_{-}$, we obtain,
$\frac{{\left(r_{+}+r_{-}-2d\right)}\left[2\Omega^{2}-\Omega^{2}_{0}\right]}{2r_{s}\sqrt{\Omega^{2}_{0}-\Omega^{2}}}-\frac{i}{\delta_{r}}\left\{\frac{\Omega}{r_{s}}\left(r_{+}(r_{+}-2d)\right.\right.\\ \left.\left.+r_{-}\left(r_{-}-2d\right)+2\left(a^{2}-k^{2}\right)\right)-2m_{l}a\right\}=-n.$ (114)
Note that the scalar’s energy levels in a rotating black hole spacetime possess hyperfine splitting, where the energy depends on the azimuthal quantum number $m_{l}$ and is directly coupled to the black hole’s angular momentum parameter $a$. This is analogous to the Zeeman effect, which occurs when a hydrogen atom is immersed in a magnetic environment. The existence of the term can be understood as an interaction between the orbiting scalar field, with magnetic state $m_{l}$, and the black hole’s angular momentum $a$.
Now, let us investigate the energy levels in the limit $(E-E_{0})r_{s}\to 0$. For modes with $\beta_{+},\gamma_{+}$ and $\beta_{-},\gamma_{-}$, we obtain,
$E_{n}=E_{0}\sqrt{1-{\left[\frac{\frac{E_{0}r_{s}}{\hbar c}\frac{{\left(r_{+}+r_{-}-2d\right)}}{r_{s}}}{2\left(i\frac{2m_{l}a}{\delta_{r}}+n\right)}\right]}^{2}},$ (115)
$E_{n}-E_{0}\approx-\frac{E_{0}}{2}{\left[\frac{\frac{E_{0}r_{s}}{\hbar c}\frac{{\left(r_{+}+r_{-}-2d\right)}}{r_{s}}}{2\left(i\frac{2m_{l}a}{\delta_{r}}+n\right)}\right]}^{2}.$ (116)
@@
$E_{n}=E_{0}\sqrt{1-{\left[\frac{\frac{E_{0}r_{s}}{\hbar c}\frac{{\left(r_{+}+r_{-}-2d\right)}}{r_{s}}}{2n}\right]}^{2}},$ (117)
$E_{n}-E_{0}\approx-\frac{E_{0}}{2}{\left[\frac{\frac{E_{0}r_{s}}{\hbar c}{\frac{{\left(r_{+}+r_{-}-2d\right)}}{r_{s}}}}{2n}\right]}^{2}.$ (118)
@@ 1. For the case with $\alpha_{+},\beta_{+},\gamma_{+}$, we obtain,
$\Omega_{2}=\frac{r_{s}\left(in(r_{+}-r_{-})+2m_{l}a\right)}{2\left(r_{-}(r_{-}-2d)+a^{2}-k^{2}\right)}.$ (120)